6. A compound of carbon and hydrogen consists of 94.34% c and 5.66% H by mass. The molecular mass of the compound is found to be 178 amu. What is its molecular formula?
Answer #1

Actual weight contributed by Carbon =(94.34/100)* 178 = 167.92 which is approx equal to 168

Actual weight contributed by H =(5.66/100)*178 =10

Therefore the molecular formula is C168/12 H 10/1 (as the atomic weight of Carbon is 12 and that of H is 1)

Thus molecular formula is C14H10